# A look at the operator product expansion in critical dynamics

### v1

@@ I Introduction
The operator product expansion (OPE) offers a non-perturbative tool to study quantum and statistical field theories. The paradigmatic example is given by two-dimensional critical phenomena, which have been studied in detail also thanks to the application of conformal field theory (CFT) techniques in which the OPE plays a key role Francesco et al. (1997). More recently, the conformal bootstrap approach allowed to obtain precise predictions also in higher dimensions Poland et al. (2019). These results have been obtained by exploiting also the powerful constraints due to conformal symmetry, which, however, may not always be present since, for instance, one may not exactly be at criticality or because of the presence of some external scale, as it happens in certain non-equilibrium systems.
@@
However, as such, the OPE holds also outside the critical region, i.e., away from a fixed point, and it can be employed to study a wide variety of systems. In particular, the OPE is used to study the expansion of the two-point function of equilibrium systems close to criticality, see, e.g., Brezin et al. (1974); Brézin et al. (1974); Caselle et al. (2015). In the context of QCD, it is employed to study deep inelastic scattering processes Novikov et al. (1978).
@@
The purpose of this paper is to study the operator product expansion in the context of critical dynamics in the case where the fluctuation-dissipation theorem applies for the first time. We shall show that the OPE allows one to define further universal quantities, besides the critical exponents, which generalize the OPE coefficients of equilibrium statistical field theories at criticality. We shall work within the response field formalism Martin et al. (1973); Janssen (1976); de Dominicis (1976), in which the so-called response field $\bar{\phi}$ is introduced over and above the physical field $\phi$. This makes the formalism very similar to what is employed in genuinely non-equilibrium systems and so it constitutes a first step towards a general study of the OPE in non-equilibrium systems. In this respect, it should be noted that a sort of short-time expansion is sometimes invoked in the studies concerning the so-called critical initial slip, which is related to initial states that are not the equilibrium ones, see Täuber (2014) and references therein. Furthermore, as time and space scale differently, our results are somehow reminiscent of the OPE for non-relativistic theories, see Golkar and Son (2014); Shimada and Shimada (2021).
@@
The paper is organized as follows. In section II we introduce our notation and conventions and briefly review some of the main features of the path integral approach to model A. In section III we introduce some composite operators of interest and analyze their renormalization and expectation value (VEV). We also briefly consider the consequences of the time-reversal symmetry at the origin of the fluctuation-dissipation theorem. In section IV we derive the Gaussian, i.e., free-field theory, OPE. We do so by considering the functional “Schroedinger” picture associated with the Fokker-Planck equation. In section V we compute the leading order corrections to the most singular parts of the OPE of the two-point and the response functions. Finally, in section VI we discuss how to extract universal information from the OPE calculated in the previous section and derive a universal form of the OPE coefficients to $O\left(\epsilon\right)$. In sections VII and VIII we consider the $O\left(N\right)$ model both perturbatively and in the large $N$ limit. We summarize our findings in section IX. Some definitions and technical discussions are confined to Appendices A and B.
@@ II.1 Model A and its path integral representation
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}\left(t,x\right)\left(\partial_{t}\phi_{0}\left(t,x\right)+D_{0}\frac{\delta H}{\delta\phi_{0}\left(t,x\right)}-D_{0}\bar{\phi}_{0}\left(t,x\right)^{2}\right)\,.$ (4)
@@ II.2 Perturbative renormalization of model A in $4-\epsilon$ dimensions
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}\left(t,x\right)\Biggr{\{}\partial_{t}\phi_{0}\left(t,x\right)+D_{0}\left(-\partial^{2}+m_{0}^{2}+\frac{g_{0}}{3!}\phi_{0}\left(t,x\right)^{2}\right)\phi_{0}\left(t,x\right)$ (6)
@@
$S$ $=$ $\int_{t,x}Z_{\bar{\phi}}^{1/2}Z_{\phi}^{1/2}\bar{\phi}\left(t,x\right)\Biggr{\{}\partial_{t}\phi\left(t,x\right)+Z_{D}D\left(-\partial^{2}+Z_{m^{2}}m^{2}+\frac{Z_{g}g}{3!}Z_{\phi}^{1/2}\phi\left(t,x\right)^{2}\right)\phi_{0}\left(t,x\right)$ (7)
$-Z_{\bar{\phi}}Z_{D}D\bar{\phi}_{0}\left(t,x\right)^{2}\Biggr{\}}\,.$
@@
Moreover, since in the static limit one recovers the equilibrium distribution, the renormalization, and so the critical exponents, of the couplings appearing in the Hamiltonian (2) are the same as in the static case. However, when discussing dynamical critical systems, a further exponent arises, the so-called dynamical critical exponent $z$, which describes the relation between spatial and temporal scalings and is related to the renormalization of the coupling $D$: $z=2+\gamma_{D}$ with $\gamma_{D}=D^{-1}\mu\frac{d}{d\mu}{D}=-Z_{D}^{-1}\mu\frac{d}{d\mu}{Z}_{D}$. At the Gaussian fixed point one has $z=2$, which implies that a two-point function takes the following scaling form: $G\left(t,x\right)=x^{-2\Delta}g\left(\frac{t}{x^{2}}\right)$.
@@ III Composite operators
In this section we consider the first order renormalization of some composite operators, which we shall use later on to construct the OPE. We also determine the normalizations of interests.
@@ III.1 The composite operator $\left[\phi^{2}/2\right]$
where actually the wave function renormalization is negligible at this order.
$Z_{20}$ $=$ $\frac{1}{16\pi^{2}\epsilon}\,.$
Figure 1: The diagram 0(a) is the diagram contributing to the VEV of $\left[\phi^{2}/2\right]$ at Gaussian order. The diagrams 0(b) contributes at order $O(g)$ while diagram 0(c) vanishes. We do not display here the diagrams associated with the counterterms stemming out of the renormalization factors $Z_{22}$, $Z_{20}$, and $Z_{m^{2}}$.
In order to have an idea about what kind of expression the VEV of $\left[\phi^{2}/2\right]$ has, let us write it down explicitly within the context of dimensional regularization:
$\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}\,=\,\Bigr{\langle}\left(\frac{\phi^{2}}{2}+\frac{m^{2}\mu^{-\epsilon}}{16\pi^{2}\epsilon}\right)\Bigr{\rangle}+O\left(g\right)$ $=$ $\frac{m^{2}}{32\pi^{2}}\left(\gamma-1+\log\frac{m^{2}}{4\pi\mu^{2}}\right)+O\left(\epsilon\right)\,,$
where we do not display explicitly the $O\left(\epsilon\right)$-terms, which however should be kept.
$Z_{ij}$ $=$ $\sum_{k}Z_{ij,k}g^{k}\,.$
At the Gaussian level, one finds $Z_{22,0}=1$ and $Z_{20,0}=\frac{1}{16\pi^{2}\epsilon}$.
To order $O\left(g\right)$ one obtains the following result for the counterterms
$Z_{22,1}$ $=$ $\frac{1}{\left(4\pi\right)^{2}}\frac{1}{\epsilon}$
$Z_{20,1}$ $=$ $\frac{1}{256\pi^{4}\epsilon^{2}}\,.$
The counterterm $Z_{22,1}$ implies the anomalous dimension $\gamma_{\phi^{2}}=-Z_{\phi^{2}}^{-1}\dot{Z}_{\phi^{2}}\approx\frac{g}{16\pi^{2}}$. At the fixed point, one obtains the following correction to the scaling dimension of the operator:
@@
where we used the fixed point value $g_{*}\approx 16\pi^{2}\epsilon/3$.
Finally, let us also consider the static limit of the two-point of $\left[\phi^{2}/2\right]$ to $O\left(g\right)$. At the fixed point, the diagrams contributing to this order are shown in Figure 3. For our purposes, we shall be interested in defining a normalization constant ${\cal N}_{2}$ such that the operator $\phi_{2}\equiv{\cal N}_{2}\left[\frac{\phi^{2}}{2}\right]$ is normalized as follows
$\Bigr{\langle}\phi_{2}\left(0,x\right)\phi_{2}\left(0,0\right)\Bigr{\rangle}\,=\,{\cal N}_{2}^{2}\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\left(0,x\right)\left[\frac{\phi^{2}}{2}\right]\left(0,0\right)\Bigr{\rangle}$ $=$ $\frac{1}{x^{2\Delta_{2}}}\,.$
After an explicit computation and performing the $\epsilon$-expansion at the fixed point, one obtains:
${\cal N}_{2}$$\approx$ $4\sqrt{2}\pi^{2}-\epsilon\frac{2}{3}\sqrt{2}\left(2\gamma\pi^{2}-\pi^{2}+2\pi^{2}\log\pi-2\pi^{2}\log\mu\right)\,.$
@@ III.2 A consequence of the time-reversal symmetry
$\phi_{0}^{\prime}\left(t,x\right)\bar{\phi}_{0}^{\prime}\left(t,x\right)$ $=$ $\phi_{0}\left(-t,x\right)\bar{\phi}_{0}\left(-t,x\right)-\frac{1}{D_{0}}\phi_{0}\left(-t,x\right)\dot{\phi}_{0}\left(-t,x\right)$
$=$ $\phi_{0}\left(-t,x\right)\bar{\phi}_{0}\left(-t,x\right)+\frac{1}{D_{0}}\partial_{t}\left(\frac{\phi_{0}\left(-t,x\right)^{2}}{2}\right)\,.$
@@
$Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi^{\prime}\left(t,x\right)\bar{\phi}^{\prime}\left(t,x\right)$ $=$ $Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(-t,x\right)\bar{\phi}\left(-t,x\right)+\frac{1}{D}\partial_{t}\left[\frac{\phi\left(-t,x\right)^{2}}{2}\right]\,.$
@@
$\Bigr{\langle}Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(t,x\right)\bar{\phi}\left(t,x\right)\Psi\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}$ $=$ $\Bigr{\langle}Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(-t,x\right)\bar{\phi}\left(-t,x\right)\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}$ (8)
$+\Bigr{\langle}\frac{1}{D}\partial_{t}\left[\frac{\phi\left(-t,x\right)^{2}}{2}\right]\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}\,,$
@@
$\left[\phi\bar{\phi}\right]$ $=$ $Z_{\phi\bar{\phi}}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(t,x\right)\bar{\phi}\left(t,x\right)$
@@ IV Gaussian operator product expansion
where $\left[\phi^{2}\right]$ and $\left[\phi\bar{\phi}\right]$ are defined in terms of the normal ordering prescription, which we shall introduce later on. Thus, the RHS of equations (10) and (11) define the OPE coefficients associated with $\phi\times\phi$ and $\phi\times\bar{\phi}$. For instance, one recognizes that the most singular coefficient of the $\phi\times\phi$ OPE is associated with the identity operator, which we shall denote by $c_{\phi\phi 1}$ with $c_{\phi\phi 1}=G_{\phi\phi}\left(t,x\right)$ in the present case. More generally, we shall denote the coefficient associated with an operator $O$ in the product $O_{1}\times O_{2}$ by $c_{O_{1}O_{2}O}$. Finally, let us note the absence in the $\phi\times\bar{\phi}$ OPE of the operator $\left[\phi^{2}\right]$ at the Gaussian level.

### v2

@@ I Introduction
The operator product expansion (OPE) Wilson (1969); Zimmermann (1973) offers a non-perturbative tool to study quantum and statistical field theories. The paradigmatic example is given by two-dimensional critical phenomena, which have been studied in detail also thanks to the application of conformal field theory (CFT) techniques in which the OPE plays a key role Francesco et al. (1997). More recently, the conformal bootstrap approach allowed to obtain precise predictions also in higher dimensions Poland et al. (2019). These results have been obtained by exploiting also the powerful constraints due to conformal symmetry, which, however, may not always be present since, for instance, one may not exactly be at criticality or because of the presence of some external scale, as it happens in certain non-equilibrium systems.
@@
However, as such, the OPE holds also outside the critical region, i.e., away from a fixed point, and it can be employed to study a wide variety of systems. In particular, the OPE is used to study the expansion of the two-point function of equilibrium systems close to criticality, see, e.g., Brezin et al. (1974); Brézin et al. (1974); Caselle et al. (2015); Costagliola (2016); Caselle et al. (2016). In the context of QCD, it is employed to study deep inelastic scattering processes Novikov et al. (1978).
@@
Given that the OPE offers such a powerful framework in the description of different physical phenomena, it is natural to investigate its application in the context of critical dynamics. The purpose of this paper is to study the operator product expansion in the context of critical dynamics in the case where the fluctuation-dissipation theorem applies for the first time. We shall work within the response field formalism Martin et al. (1973); Janssen (1976); de Dominicis (1976), in which the so-called response field $\bar{\phi}$ is introduced over and above the physical field $\phi$. This makes the formalism very similar to what is employed in genuinely non-equilibrium systems and so it constitutes a first step towards a general study of the OPE in non-equilibrium systems. In the context of critical dynamics, one of the main motivations behind the present work is indeed to perform some basic analysis and some concrete calculations of OPE coefficients, which has been far less explored than its static counterpart. As such, this work has an exploratory field theoretic character. Let us note that a sort of short-time expansion is sometimes invoked in the studies concerning the so-called critical initial slip, which is related to initial states that are not the equilibrium ones, see Täuber (2014) and references therein. Furthermore, as time and space scale differently, our results are somehow reminiscent of the OPE for non-relativistic theories, see Golkar and Son (2014); Shimada and Shimada (2021). Finally, let us anticipate that, besides the field theoretic aspects, we shall show that the OPE allows one to define further universal quantities, besides the critical exponents, which generalize the OPE coefficients of equilibrium statistical field theories at criticality.
@@
The paper is organized as follows. In section II we introduce our notation and conventions and briefly review some of the main features of the path integral approach to model A. In section III we introduce some composite operators of interest and analyze their renormalization and expectation value (VEV). We also briefly consider the consequences of the time-reversal symmetry at the origin of the fluctuation-dissipation theorem. In section IV we derive the Gaussian, i.e., free-field theory, OPE. We do so by considering the functional “Schroedinger” picture associated with the Fokker-Planck equation. In section V we compute the leading order corrections to the most singular parts of the OPE of the two-point and the response functions. Finally, in section VI we discuss how to extract universal information from the OPE calculated in the previous section and derive a universal form of the OPE coefficients to $O\left(\epsilon\right)$. In sections VII and VIII we consider the $O\left(N\right)$ model both perturbatively and in the large $N$ limit. We summarize our findings in section IX. Some definitions and technical discussions are confined to several Appendices. We highlight graphically the main results of our work by encircling the most relevant expressions in a box.
@@ II.1 Model A and its path integral representation
$S$ $=$ $\int_{t,x}\left[\bar{\phi}_{0}\left(t,x\right)\left(\partial_{t}\phi_{0}\left(t,x\right)+D_{0}\frac{\delta H}{\delta\phi_{0}\left(t,x\right)}\right)-D_{0}\bar{\phi}_{0}\left(t,x\right)^{2}\right]\,.$ (4)
@@ II.2 Perturbative renormalization of model A in $4-\epsilon$ dimensions
$S$ $=$ $\int_{t,x}\Biggr{\{}\bar{\phi}_{0}\left(t,x\right)\Biggr{[}\partial_{t}\phi_{0}\left(t,x\right)+D_{0}\left(-\partial^{2}+m_{0}^{2}+\frac{g_{0}}{3!}\phi_{0}\left(t,x\right)^{2}\right)\phi_{0}\left(t,x\right)\Biggr{]}$ (6)
@@
$S$ $=$ $\int_{t,x}\Biggr{\{}Z_{\bar{\phi}}^{1/2}Z_{\phi}^{1/2}\bar{\phi}\left(t,x\right)\Biggr{[}\partial_{t}\phi\left(t,x\right)+Z_{D}D\left(-\partial^{2}+Z_{m^{2}}m^{2}+\frac{Z_{g}g}{3!}Z_{\phi}^{1/2}\phi\left(t,x\right)^{2}\right)\phi\left(t,x\right)\Biggr{]}$ (7)
$-Z_{\bar{\phi}}Z_{D}D\bar{\phi}\left(t,x\right)^{2}\Biggr{\}}\,.$
@@
Moreover, since in the static limit one recovers the equilibrium distribution, the renormalization, and so the assiociated critical exponents, of the couplings appearing in the Hamiltonian (2) are the same as in the static case. However, when discussing dynamical critical systems, a further exponent arises, the so-called dynamical critical exponent $z$, which describes the relation between spatial and temporal scalings and is related to the renormalization of the coupling $D$: $z=2+\gamma_{D}$ with $\gamma_{D}=D^{-1}\mu\frac{d}{d\mu}{D}=-Z_{D}^{-1}\mu\frac{d}{d\mu}{Z}_{D}$. At the Gaussian fixed point one has $z=2$, which implies that a two-point function takes the following scaling form: $G\left(t,x\right)=x^{-2\Delta}g\left(\frac{t}{x^{2}}\right)$.
@@ III Composite operators
The operator product expansion allows one to express the insertion of two composite operators in a correlation function in the short distance limit as a sum of insertions of a single composite operator multiplied by suitable coefficients, the so-called Wilson coefficients Wilson (1969); Zimmermann (1973). In practice, the OPE states the validity of
$\left[O_{i}\left(x\right)\right]\left[O_{j}\left(y\right)\right]=\sum_{k}C_{ijk}\left(x-y\right)\left[O_{k}\left(\frac{x+y}{2}\right)\right]\,,$ (8)
inserted into any correlation function in the limit of $\left|x-y\right|$ approaching zero. In (8), $C_{ijk}\left(x-y\right)$ indicates a Wilson coefficient and the operators $\left[O_{i}\right]\left[O_{j}\right]$ are taken to be time ordered.
It must be emphasized that the composite operators $\left[O_{i}\right]$ are defined as renormalized composite operators, i.e., they are not mere field monomials, such as $\phi^{2}$, but they are modified by suitable renormalization factors. In this section, within perturbation theory, we consider the first order renormalization of certain composite operators, which we shall use later on to compute the most singular parts of the OPE of the two-point function and the response function. In the case of the two-point function, one has $\left[O_{i}\right]=\left[O_{j}\right]=\phi$ while in the case of the response function $\left[O_{i}\right]=\phi$ and $\left[O_{j}\right]=\bar{\phi}$. We also determine certain normalizations factors of interests.
@@ III.1 The composite operator $\left[\phi^{2}/2\right]$
where actually the wave function renormalization is negligible at this order. We shall limit ourselves to list some results and refer the redear to Appendix B for more details.
We begin our discussion by considering the VEV of $\left[\phi^{2}/2\right]$ and its expression at tree level. Let us write it down explicitly within the context of dimensional regularization:
$\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}\,=\,\Bigr{\langle}\left(\frac{\phi^{2}}{2}+\frac{m^{2}\mu^{-\epsilon}}{16\pi^{2}\epsilon}\right)\Bigr{\rangle}+O\left(g\right)$ $=$ $\frac{m^{2}}{32\pi^{2}}\left(\gamma-1+\log\frac{m^{2}}{4\pi\mu^{2}}\right)+O\left(\epsilon\right)\,,$ (9)
where we do not display explicitly the $O\left(\epsilon\right)$-terms, which however should be kept. As far as the scaling dimension of the operator is concerned, one obtains the following expression to the first order in the $\epsilon$-expansion:
@@
Finally, let us also consider the static limit of the two-point of $\left[\phi^{2}/2\right]$. For our purposes, we shall be interested in defining a normalization constant ${\cal N}_{2}$ such that the operator $\phi_{2}\equiv{\cal N}_{2}\left[\frac{\phi^{2}}{2}\right]$ is normalized as follows
$\Bigr{\langle}\phi_{2}\left(0,x\right)\phi_{2}\left(0,0\right)\Bigr{\rangle}\,=\,{\cal N}_{2}^{2}\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\left(0,x\right)\left[\frac{\phi^{2}}{2}\right]\left(0,0\right)\Bigr{\rangle}$ $=$ $\frac{1}{x^{2\Delta_{2}}}\,.$ (11)
The explicit value of ${\cal N}_{2}$ is given in Appendix B.
@@ III.2 A consequence of the time-reversal symmetry
$\phi_{0}^{\prime}\left(t,x\right)\bar{\phi}_{0}^{\prime}\left(t,x\right)$ $=$ $\phi_{0}\left(-t,x\right)\bar{\phi}_{0}\left(-t,x\right)-\frac{1}{D_{0}}\phi_{0}\left(-t,x\right)\dot{\phi}_{0}\left(-t,x\right)$ (12)
$=$ $\phi_{0}\left(-t,x\right)\bar{\phi}_{0}\left(-t,x\right)+\frac{1}{D_{0}}\partial_{t}\left(\frac{\phi_{0}\left(-t,x\right)^{2}}{2}\right)\,.$ (13)
@@
$Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi^{\prime}\left(t,x\right)\bar{\phi}^{\prime}\left(t,x\right)$ $=$ $Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(-t,x\right)\bar{\phi}\left(-t,x\right)+\frac{1}{D}\partial_{t}\left[\frac{\phi\left(-t,x\right)^{2}}{2}\right]\,.$ (14)
@@
$\Bigr{\langle}Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(t,x\right)\bar{\phi}\left(t,x\right)\Psi\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}=$ (15)
$\Bigr{\langle}Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(-t,x\right)\bar{\phi}\left(-t,x\right)\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}+\Bigr{\langle}\frac{1}{D}\partial_{t}\left[\frac{\phi\left(-t,x\right)^{2}}{2}\right]\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}\,,$
@@
$\left[\phi\bar{\phi}\right]$ $=$ $Z_{\phi\bar{\phi}}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(t,x\right)\bar{\phi}\left(t,x\right)$ (17)
@@ IV Gaussian operator product expansion
where $\left[\phi^{2}\right]$ and $\left[\phi\bar{\phi}\right]$ are defined in terms of the normal ordering prescription, which we shall introduce later on. The perturbative corrections to (20) and (21) will be calculated in section V. Let us note that in the interacting theory the normal ordering prescription will no longer be sufficient to define a composite operator and we will consider the composite operators introduced in section III. In this section, however, we find it useful to derive the Gaussian OPE by adopting the normal ordering prescription.
The RHS of equations (20) and (21) define the OPE coefficients associated with $\phi\times\phi$ and $\phi\times\bar{\phi}$. For instance, one recognizes that the most singular coefficient of the $\phi\times\phi$ OPE is associated with the identity operator, which we shall denote by $C_{\phi\phi 1}$ with $C_{\phi\phi 1}=G_{\phi\phi}\left(t,x\right)$ in the present case. More generally, we shall denote the coefficient associated with an operator $O$ in the product $O_{1}\times O_{2}$ by $C_{O_{1}O_{2}O}$. Finally, let us note the absence in the $\phi\times\bar{\phi}$ OPE of the operator $\left[\phi^{2}\right]$ at the Gaussian level.